OpenStudy (big_mensa961):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (welshfella):

the general form is (x - a)^2 + (y - b)^2 = r^2 where (a,b) is the center and r = radius

OpenStudy (big_mensa961):

is this right\[(x-5)^2+(y-6)^2=3\]

OpenStudy (greenbannanabob):

The equation of a circle is (x – h)^2 + (y – k)^2 = r^2 The radius is r The x value of the center is h and the y value is k

OpenStudy (greenbannanabob):

3 squared at the end

OpenStudy (welshfella):

its not 3 its 3^2

OpenStudy (big_mensa961):

\[(x−5)2+(y−6)2=3^2\]

OpenStudy (welshfella):

otherwise ok

OpenStudy (greenbannanabob):

That's correct